ssignment Saved A financial planning service offers a college savings program. The plan calls for you to ach, with the first payment occurring today, your childs 12th birthday. Beginning on your childs 18th birthday, the plan will provide $35,000 per year for four years. What return is this investment offering? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Return
Answer #1

Total Program Payments to be made in 6 years = $14,800 x 6 = $88,800

Starting 7th year (Child’s 18th Birthday) Returns for 4 Years from the Program = 35,000 x 4 = $140,000

Since, Last/4th Payment (of $35,000) made occurs at the end of 10th year

Total Payments received in 10 Years = $140,000

Investment made in 10 years (ignoring PV/FV Factors) = $88,800

Net Return in 10 Years would be $140,000 – 88,800 = $51,200

= 51,200/88,800 /10 = 0.0576 = 5.76%